OpenStudy (anonymous):

Solve the inequality (2x-7)/(x-5) greater than/equal to 3

OpenStudy (gudden):

let in the first case it be equal to 3 (2x-7)/(x-5) = 3 it means (2x-7) = 3 (x-5) solving it gives 2x - 7 = 3x - 15 2x - 3x = -15 + 7 - x = - 8 which gives x = 8 and let in the second case (2x-7)/(x-5) > 3 solving it further gives (2x-7) > (x-5) 3 further solving it gives 2x - 7 > 3x - 15 now try solving it further... and if any problem comes just ask me... i would feel very happy in clearing ur concept

OpenStudy (anonymous):

Would you solve it this way any time there is a fraction with a number on the other side of the inequality?

OpenStudy (anonymous):

Also, would the horizontal asymptote be y=0?

OpenStudy (gudden):

ofcourse.. you would solve it that ways.. @kaylady21 and yes.. y = 0..

OpenStudy (ybarrap):

(2x-7)/(x-5) >= 3 2x-7 >= 3(x-5) 2x-7 >= 3x-15) -7 + 15 >= 3x -2x = x 8 >= x x <= 8